# The HST-Hyperion Survey: Environmental Imprints on the Stellar-Mass Function at $z\sim 2.5$

### v1

@@ 2.3 HST Slitless Spectroscopy
In addition to the COSMOS2020 photometry and ground-based spectroscopy, we incorporate HST slitless (grism) spectroscopy from the HST-Hyperion Survey to greatly increase the number of secure redshifts in our sample. While details regarding target selection, data reduction, and redshift classification are provided in Forrest et al. (2025), we briefly outline the relevant information here. HST-Hyperion was an HST Cycle 29 program (PI: Lemaux) which obtained 50 orbits of WFC3/G141 slitless spectroscopy with WFC3/F160W imaging in 25 pointings covering the densest peaks in Hyperion. Additionally, the survey included a re-analysis of all 3D-HST data in the field as well (Brammer et al. 2012; Momcheva et al. 2016). In total, the observations resulted in $\sim 12,800$ spectra for objects with $m_{\rm NIR}<25$.
@@
Following this process, a final quality flag was assigned to each object, ranging from 0 (lowest quality) to 5 (highest quality). To calibrate the reliability of the quality flags, we compared the HST-grism redshift ($z_{g}$) with the ground-based spectroscopic redshift ($z_{s}$) for objects with both measurements. We only considered grism spectra with flags $\geq 3$ and ground-based spectra with quality flags of 3 or 4 (see Section 2.2.1), treating the highly reliable ground-based redshifts as the true values. We then examined the distribution of $\Delta z/(1+z_{s})\equiv(z_{g}-z_{s})/(1+z_{s})$ and calculated the normalized median absolute deviation ($\sigma_{\rm NMAD}$; Hoaglin et al. 1983), a measure of the distribution’s width which is more robust to outliers. Denoting the median of a value $x$ as $\text{Med}(x)$, $\sigma_{\rm NMAD}$ is given as
@@
For the combined samples of grism redshifts with quality flags $\geq 3$, we find $\sigma_{\rm NMAD}\sim 0.0016$, corresponding to a median wavelength offset of $\sim 22\,\text{\AA }$ at $z=2.5$. Using a reliability threshold of $\Delta z/(1+z_{s})<0.004$, we determined the fraction of spectra within this cutoff for each quality flag. The reliability for flags 3, 4, and 5 is 56.7%, 77.3%, and 90.4%, respectively. In total, we have grism redshifts for 5,338 objects which have reliability $\geq 56.7\%$, have associated COSMOS2020 photometry, are identified as either galaxies or X-ray sources in COSMOS, lay in the ranges $\rm 149.60^{o}\leq RA\leq 150.52^{o}$ and $\rm 1.74^{o}\leq Dec\leq 2.73^{o}$, and have either $3.6]\leq 24.8$ or $4.5]\leq 24.8$ (per COSMOS2020). The redshift distribution of these sources is shown in the bottom panel of Figure 2.
@@ 2.4 An Updated View of Hyperion
Hyperion is a massive, overdense structure centered at $z\sim 2.5$, exhibiting multiple density peaks linked by filamentary bridges. Because such “protostructures” are not yet virialized, their overdensity field is highly inhomogeneous, unlike the smoother, monotonic profiles of relaxed $z\!\lesssim\!0.5$ clusters. Throughout this paper we therefore describe Hyperion in overdensity space: the highest‐density peaks are wrapped in moderately overdense outskirts, beyond which lies the lower-density field. Formal overdensity thresholds for the peaks, outskirts, and field are given in Section 3.2.
@@
In this study, we use an updated VMC map as the reference for the underlying structure of Hyperion (see Section 3 for a brief summary, or a complete description in Lemaux et al. 2022). Leveraging the updated photometric catalog and improved spectroscopic completeness provided by the new HST data described above, we are able to populate Hyperion with more member galaxies and provide an updated view of the Hyperion proto-supercluster in Figure 11. As mapped with the VMC technique, Hyperion lies in the ranges $\rm 149.88^{o}\leq RA\leq 150.41^{o}$, $\rm 2.03^{o}\leq decl.\leq 2.50^{o}$, and $2.41\leq z\leq 2.68$. The total mass is estimated to be $\log_{10}(M/M_{\odot})\sim 15.40$ and the total volume is estimated to be $\rm\sim 4.3\times 10^{4}\,\,cMpc$. While the most overdense region of Hyperion lays at $z\lesssim 2.52$, a high-redshift “offshoot” appears to extend up until $z\sim 2.68$. A more detailed discussion of the structure of Hyperion, its constituent peaks, and the inclusion of the high-redshift offshoot is provided in Appendix A.
@@ 3 Characterizing Galaxy Environments
Developing a quantitative method to describe the environment in which a galaxy resides is central to understanding the environmental effects on galaxy evolution. While there are a variety of methods used to quantify local environments, we employed a Voronoi Monte Carlo mapping technique which has proven successful in many previous works (e.g., Lemaux et al. 2017; Cucciati et al. 2018; Hung et al. 2020; Shen et al. 2021; Lemaux et al. 2022; Forrest et al. 2023; Staab et al. 2024). This method produces a 3D grid of voxels, each with an associated overdensity value, to which galaxies are assigned. From these maps, we can identify extended, connected regions of voxels exceeding a given overdensity threshold, which are taken to be protostructures. In this section, we briefly explain the Voronoi Monte Carlo mapping technique before discussing the specific details of the map used in this study.
@@ 3.1.1 The Voronoi Monte Carlo Method
The Voronoi Monte Carlo (VMC) algorithm leverages a combination of well-constrained spectroscopic redshifts and less-constrained photometric redshifts to statistically map the 3D spatial distribution of galaxies. Broadly speaking, a Voronoi tessellation map divides a plane into polygons, where each polygon is centered on a specific point in a given distribution. These polygons encompass all locations in the plane that are closer to their corresponding point than to any other point in the distribution, effectively partitioning the plane based on proximity. In the context of mapping galaxy density fields, the polygons partition two dimensions (R.A. and decl.) based on proximity to the closest galaxy (see Figure 3 of Tomczak et al. 2017).
Given the large pecular velocities of galaxie within protostructures, the size of the uncertainties for the photometric redshifts, and the sparseness of the spectroscopic sample, this method cannot be directly applied in the redshift dimension. Instead, the tessellation maps are constructed in the 2D R.A.-decl. plane in overlapping redshift slices. A set of Monte Carlo realizations is generated for each redshift slice of the 3D space in order to minimize the effects of photometric redshift uncertainties by better sampling the redshift PDF. In a given realization, each galaxy in the data set is assigned a redshift based on the following logic:
  1. 1.
    Galaxies without spectroscopic redshift: A new redshift is drawn from an asymmetric Gaussian distribution centered on the reported photometric redshift and with lower/upper standard deviations equal to the lower/upper uncertainties ascribed to their redshift probability distributions.
  2. 2.
    Galaxies with spectroscopic redshift: Each quality flag has an associated likelihood and uncertainty (e.g. $0.70_{-0.06}^{+0.04}$ for flags X2/X9), and threshold for the realization is drawn from an asymmetric Gaussian with the associated mean and dispersion. A random number is also drawn from a uniform distribution ranging from 0 to 100. If the number is less than or equal to the threshold associated with the quality-flag of the spectroscopic redshift (e.g. $\lesssim 70_{-6}^{+4}$ for flags X2/X9), the reported spectroscopic redshift is kept. Otherwise, a new redshift is drawn from the photometric redshift distribution, as described above.
Only the galaxies assigned redshifts in the given redshift slice are kept for each realization, and a 2D Voronoi tessellation is performed to create a 2D polygon map. The density associated with a polygon, $\Sigma_{\alpha,\delta,z}$, is then inversely proportional to its area in $\rm Mpc^{2}$. This is re-mapped onto a consistent grid of voxels in R.A.-decl. space separated by 75 pkpc, where each voxel has the density of the polygon enclosing its center in the given Monte Carlo realization.
@@ 3.1.2 Applying the VMC Method
The VMC maps used in the present study are described in Lemaux et al. (2022), and thus we give only a brief description here. The maps were generated using the COSMOS2015 photometry (Laigle et al. 2016) and associated field spectroscopic surveys (described in Section 2.2.1). Targeted spectroscopic surveys were used primarily to improve statistics at redshifts higher than Hyperion in and effort to mitigate biases in overdensities of targeted, intermediate-redshift protostructures. The VMC process, as described above, was performed over the entire COSMOS footprint for redshifts $2\leq z\leq 5$. The redshift bins had widths of $7.5\,\,\text{Mpc}$, with overlaps of 90% between adjacent redshift slices. A total of 100 MC realizations were performed for each redshift slice to generate the map, using only those galaxies that fell within the redshift bin and with magnitudes $[3.6]<24.8$ for the density calculations in a given realization. Afterwards, each voxel was assigned a density $\Sigma_{\alpha,\delta,z}$ equal to the median density of that voxel across all MC realizations.
@@
In order to find $\langle\Sigma_{z}\rangle$, the outer 10% of voxels in each redshift slice are first trimmed to mitigate edge effects. The remaining 2D density distribution in the redshift slice, $\Sigma_{z}$, is then smoothed with a 2D-Gaussian kernel with standard deviations of three pixels in the two transverse dimensions (R.A. and decl.). Additionally, a boxcar smooth is performed in the redshift dimension with a width of three slices. Since $\Sigma_{z}$ roughly follows a log-normal distribution, a Gaussian is fit to the distribution of $\log_{10}(\Sigma_{z})$ in each redshift slice. Finally, a $2\sigma$-clipping is applied and another Gaussian is fit, resulting in a mean $\mu_{\Sigma}(z)$ and standard deviation $\sigma_{\Sigma}(z)$ for each redshift slice (top left panel of Figure 3). The mean density in the redshift-bin is related to the mean and standard deviation of the log-normal distribution via $\langle\Sigma_{z}\rangle=10^{\mu}\cdot\exp(2.652\sigma^{2})$.
Once the average density is found in each redshift slice, the overdensity can be calculated for each voxel in the slice from Equation 2. As above, the distribution of $\log(1+\delta_{\rm gal})$ in the slice is fit with a $2\sigma$-clipped Gaussian resulting in a $\mu_{\delta}(z)$ and $\sigma_{\delta}(z)$ for that slice (bottom left panel of Figure 3). It should be noted that, given that protostructures are extended in redshift space over multiple redshift slices, it is possible that an entire redshift slice is over- or underdense. Thus, a 5th-order polynomial is fit to $\mu_{\delta,z}$ and $\sigma_{\delta,z}$ to better characterize the general variation across redshift (see right panels of Figure 3), resulting in measurements for the parameters as a function of redshift, $\mu_{\delta,5}(z)$ and $\sigma_{\delta,5}(z)$.
With these in hand, we now have a metric for characterizing the local overdensity in each voxel of the VMC map by representing it as:
@@
Figure 3: Overdensity Mapping – Left: The distributions (with median set to zero) of log-density (top), and local-overdensity (bottom) for all of the voxels in the redshift bin centered on $z\sim 2.51$. The Gaussian and $2\sigma$-clipped Gaussian (red and green, respectively) are fit over both distributions. Right: The distribution of the mean-overdensity (top) and standard deviation of overdensity (bottom) as a function of redshift. The functions are fit with a $\rm 5^{th}$-order polynomial (green line) in order to smooth over small fluctuations in redshift-space.
@@ 3.2 Identifying Protostructures
In the VMC overdensity maps we call a region a protostructure if it satisfies all three of the following criteria:
@@
    it is a contiguous set of voxels with overdensity $n_{\sigma}\geq 2.5$;
@@
    it contains at least one voxel with $n_{\sigma}\geq 4.0$ (i.e. an overdensity peak);
@@
    the voxels together enclose a total mass $M_{\rm tot}\geq 10^{13}\,M_{\odot}$.
In our maps, $n_{\sigma}\geq 2.5$ and $n_{\sigma}\geq 4.0$ correspond to $\log_{10}(1+\delta_{\rm gal})\sim 0.21$ and $\log_{10}(1+\delta_{\rm gal})\sim 0.35$ at $z\sim 2.5$, respectively. The mass of the voxels is estimated by using (from Steidel et al. 1998):
@@
where $\bar{\rho}$ is the mean comoving density of the universe, $\langle\delta_{\rm gal}\rangle$ is the mean overdensity of the voxels defining the peak, $V$ is the combined comoving volume of the voxels, and $b$ is the bias factor, (2.55, as derived at $z\sim 2.5$ by Durkalec et al. 2015). It should be noted that, while this mass estimate may be rough, in this work it is used primarily as a method for differentiating between regions that are overdense because they are potentially protostructures, and regions that are overdense simply due to cosmological perturbations in density.
With this framework established, Hyperion is defined as the most massive structure in the COSMOS field. Furthermore, by considering only voxels within Hyperion above higher $n_{\sigma}$-thresholds, we can isolate its overdensity peaks in our analysis. Based on criteria (ii) of the protostructure definition, we define the peaks of Hyperion to be contiguous sets of voxels within Hyperion which have $n_{\sigma}\geq 4.0$. The outskirts of Hyperion are then the remaining voxels which fall in the overdensity range $2.5\leq n_{\sigma}\leq 4.0$. A more detailed description of Hyperion is given in Appendix A.

### v2

@@ 2.3 HST slitless spectroscopy
In addition to the COSMOS2020 photometry and ground-based spectroscopy, we incorporated HST slitless (grism) spectroscopy from the HST-Hyperion Survey to greatly increase the number of secure redshifts in our sample. While details regarding target selection, data reduction, and redshift classification are provided in Forrest et al. (2025), we briefly outline the relevant information here. HST-Hyperion was an HST Cycle 29 program (PI: Lemaux) which obtained 50 orbits of WFC3/G141 slitless spectroscopy with WFC3/F160W imaging in 25 pointings covering the densest peaks in Hyperion. Additionally, the survey included a reanalysis of all 3D-HST data in the field as well (Brammer et al. 2012; Momcheva et al. 2016). In total, the observations resulted in $\sim 12,800$ spectra for objects with $m_{\rm NIR}<25$.
@@
Following this process, a final quality flag was assigned to each object, ranging from 0 (lowest quality) to 5 (highest quality). To calibrate the reliability of the quality flags, we compared the HST-grism redshift ($z_{g}$) with the ground-based spectroscopic redshift ($z_{s}$) for objects with both measurements. We only considered grism spectra with flags $\geq 3$ and ground-based spectra with quality flags of 3 or 4 (see Section 2.2.1), treating the highly reliable ground-based redshifts as the true values. We then examined the distribution of $\Delta z/(1+z_{s})\equiv(z_{g}-z_{s})/(1+z_{s})$ and calculated the normalized median absolute deviation ($\sigma_{\rm NMAD}$; Hoaglin+83), a measure of the distribution’s width that is more robust to outliers. Denoting the median of a value $x$ as $\text{Med}(x)$, $\sigma_{\rm NMAD}$ is given as
@@
For the combined samples of grism redshifts with quality flags $\geq 3$, we find $\sigma_{\rm NMAD}\sim 0.0016$, corresponding to a median wavelength offset of $\sim 22\,\text{\AA }$ at $z=2.5$. Using a reliability threshold of $\Delta z/(1+z_{s})<0.004$, we determined the fraction of spectra within this cutoff for each quality flag. The reliability for flags 3, 4, and 5 is 56.7%, 77.3%, and 90.4%, respectively. In total, we have grism redshifts for 5,338 objects which have reliability $\geq 56.7\%$, have associated COSMOS2020 photometry, are identified as either galaxies or X-ray sources in COSMOS, lie in the ranges $\rm 149.60^{o}\leq R.A.\leq 150.52^{o}$ and $\rm 1.74^{o}\leq Dec.\leq 2.73^{o}$, and have either $3.6]\leq 24.8$ or $4.5]\leq 24.8$ (per COSMOS2020). The redshift distribution of these sources is shown in the bottom panel of Figure 2.
@@ 2.4 An updated view of Hyperion
Hyperion is a massive, overdense structure centered at $z\sim 2.5$, exhibiting multiple density peaks linked by filamentary bridges, as seen in Figure 3. Because such protostructures are not yet virialized, their overdensity field is highly inhomogeneous, unlike the smoother, monotonic profiles of relaxed $z\!\lesssim\!0.5$ clusters. Throughout this paper we therefore describe Hyperion in overdensity space: the highest‐density peaks (green in Figure 3) are wrapped in moderately overdense outskirts (purple in Figure 3), beyond which lies the lower-density field. Formal overdensity thresholds for the peaks, outskirts, and field are given in Section 3.2.
@@
In this study, we used an updated overdensity map as the reference for the underlying structure of Hyperion (see Section 3 for a brief summary, or a complete description in Lemaux et al. 2022). As mapped with the VMC method, Hyperion lies in the ranges $\rm 149.88^{o}\leq R.A.\leq 150.41^{o}$, $\rm 2.03^{o}\leq Dec.\leq 2.50^{o}$, and $2.41\leq z\leq 2.68$ (shown in Figure 3). The total mass is estimated to be $\log_{10}(M/M_{\odot})\sim 15.40$ and the total volume is estimated to be $\rm\sim 4.3\times 10^{4}\,\,cMpc$. While the most overdense region of Hyperion lies at $z\lesssim 2.52$, a high-redshift offshoot appears to extend up until $z\sim 2.68$. A more detailed discussion of the structure of Hyperion, its constituent peaks, and the inclusion of the high-redshift offshoot is provided in Appendix A.
Leveraging the updated photometric catalog and improved spectroscopic completeness provided by the new HST data described above, we were able to populate Hyperion with more member galaxies. The approximate locations of these members are shown in Figure 10, although much of the following analysis relies on the full redshift probability distribution functions for each galaxy (see Section 4.1) rather than the discrete redshift estimates plotted in Figure 10. While the relative fractions of star-forming and quiescent galaxies in Hyperion are beyond the scope of this work and will be explored in a forthcoming analysis, it is well established that the vast majority of galaxies at $z\sim 2.5$ are star-forming (Weaver et al. 2023). However, quenched galaxies – though expected to be rare at this epoch – were not systematically excluded from our analysis. Because quenched systems at these redshifts are typically massive ($\log_{10}(M_{*}/M_{\odot})\gtrsim 10.2$ Weaver et al. 2023), their redshifts can be constrained via continuum emission in the HST/grism data even in the absence of strong emission lines (Forrest et al. 2025). We therefore expect that, if a small population of quiescent galaxies is present in Hyperion, it is certainly represented in our sample in some capacity.
Figure 3: Left: 3D rendering of Hyperion, as defined by the VMC technique. Right: 2D projections of Hyperion in the (top) R.A.-Dec., (middle) R.A.-$z$, and (bottom) Dec.-$z$ planes. In all panels, the contours delineate regions of differing overdensity within the VMC map. The purple and green shaded regions correspond to Hyperion’s outskirts ($\rm 2.5\leq n_{\sigma}\leq 4.0$) and peaks ($\rm n_{\sigma}\geq 4.0$), respectively, where $\rm n_{\sigma}$ is defined in Equation 2. Also shown in red are regions with $\rm n_{\sigma}\geq 5.0$, which generally trace the most massive peaks of Hyperion (detailed in Table 2).
@@ 3 Characterizing galaxy environments
Developing a quantitative method to describe the environment in which a galaxy resides is central to understanding the environmental effects on galaxy evolution. While there are a variety of methods used to quantify local environments, we employed a VMC mapping technique which has proven successful in many previous works (e.g., Lemaux et al. 2017; Cucciati et al. 2018; Hung et al. 2020; Shen et al. 2021; Lemaux et al. 2022; Forrest et al. 2023; Staab et al. 2024). This method produces a 3D grid of voxels, each with an associated overdensity value, to which galaxies are assigned. From these maps, we identified extended, connected regions of voxels exceeding a given overdensity threshold, which were taken to be protostructures. In this section, we briefly explain the VMC mapping technique before discussing the specific details of the map used in this study.
We note that the selection criteria imposed on the galaxy populations used in both the VMC mapping and the SMF analysis are largely consistent. In particular, no explicit color cut is applied, and the only magnitude selection is based on the IRAC magnitudes of the objects ($3.6]\leq 24.8$ or $4.5]\leq 24.8$), which correspond to rest-frame near-infrared emission at $z\sim 2.5$ and therefore provide a tracer of stellar mass for Hyperion members. As discussed in Section 2.4, the galaxy population at these redshifts is expected to be dominated by star-forming systems (Weaver et al. 2023). Consequently, although no explicit selection on star formation activity is imposed, the overdensity map generated via the VMC method primarily traces the spatial distribution of star-forming galaxies due to their substantially higher number density.
@@ 3.1.1 The Voronoi Monte Carlo method
The VMC algorithm leverages a combination of well-constrained spectroscopic redshifts and less-constrained photometric redshifts to statistically map the 3D spatial distribution of galaxies. Broadly speaking, a Voronoi tessellation map divides a plane into polygons, where each polygon is centered on a specific point in a given distribution. These polygons encompass all locations in the plane that are closer to their corresponding point than to any other point in the distribution, effectively partitioning the plane based on proximity. In the context of mapping galaxy density fields, the polygons partition two dimensions (R.A. and Dec.) based on proximity to the closest galaxy (see Figure 3 of Tomczak et al. 2017).
Given the large peculiar velocities of galaxies within protostructures, the size of the uncertainties for the photometric redshifts, and the sparseness of the spectroscopic sample, this method cannot be directly applied in the redshift dimension. Instead, the tessellation maps are constructed in the 2D R.A.-Dec. plane in overlapping redshift slices. A set of Monte Carlo realizations is generated for each redshift slice of the 3D space in order to minimize the effects of photometric redshift uncertainties by better sampling the redshift PDF. In a given realization, each galaxy in the dataset is assigned a redshift based on the following logic:
Galaxies without spectroscopic redshift:
A new redshift is drawn from an asymmetric Gaussian distribution centered on the reported photometric redshift and with lower (upper) standard deviations equal to the lower (upper) uncertainties ascribed to their redshift probability distributions.
Galaxies with spectroscopic redshift:
Each quality flag has an associated likelihood and uncertainty (e.g., $0.70_{-0.06}^{+0.04}$ for flags X2/X9), and threshold for the realization is drawn from an asymmetric Gaussian with the associated mean and dispersion. A random number is also drawn from a uniform distribution ranging from 0 to 100. If the number is less than or equal to the threshold associated with the quality-flag of the spectroscopic redshift (e.g., $\lesssim 70_{-6}^{+4}$ for flags X2/X9), the reported spectroscopic redshift is kept. Otherwise, a new redshift is drawn from the photometric redshift distribution, as described above.
Only the galaxies assigned redshifts in the given redshift slice are kept for each realization, and a 2D Voronoi tessellation is performed to create a 2D polygon map. The density associated with a polygon, $\Sigma_{\alpha,\delta,z}$, is then inversely proportional to its area in $\rm Mpc^{2}$. This is remapped onto a consistent grid of voxels in R.A.-Dec. space separated by 75 pkpc, where each voxel has the density of the polygon enclosing its center in the given Monte Carlo realization.
@@ 3.1.2 Applying the VMC method
The VMC maps used in the present study are described in Lemaux et al. (2022), and thus we give only a brief description here. The maps were generated using the COSMOS2015 photometry (Laigle et al. 2016) and associated field spectroscopic surveys (described in Section 2.2.1). Targeted spectroscopic surveys were used primarily to improve statistics at redshifts higher than Hyperion in an effort to mitigate biases in overdensities of targeted, intermediate-redshift protostructures. The VMC process, as described above, was performed over the entire COSMOS footprint for redshifts $2\leq z\leq 5$. The redshift bins had widths of $7.5\,\,\text{Mpc}$, with overlaps of 90% between adjacent redshift slices. A total of 100 MC realizations were performed for each redshift slice to generate the map, using only those galaxies that fell within the redshift bin and with magnitudes $[3.6]<24.8$ for the density calculations in a given realization. Afterwards, each voxel was assigned a density $\Sigma_{\alpha,\delta,z}$ equal to the median density of that voxel across all MC realizations.
@@
In order to find $\langle\Sigma_{z}\rangle$, the outer 10% of voxels in each redshift slice were first trimmed to mitigate edge effects. The remaining 2D density distribution in the redshift slice, $\Sigma_{z}$, was then smoothed with a 2D-Gaussian kernel with standard deviations of three pixels in the two transverse dimensions (R.A. and Dec.). Additionally, a boxcar smooth was performed in the redshift dimension with a width of three slices. Since $\Sigma_{z}$ roughly follows a log-normal distribution, a Gaussian was fit to the distribution of $\log_{10}(\Sigma_{z})$ in each redshift slice. Finally, a $2\sigma$-clipping was applied and another Gaussian was fit, resulting in a mean $\mu_{\Sigma}(z)$ and standard deviation $\sigma_{\Sigma}(z)$ for each redshift slice (top left panel of Figure 4). The mean density in the redshift-bin is related to the mean and standard deviation of the log-normal distribution via $\langle\Sigma_{z}\rangle=10^{\mu}\cdot\exp(2.652\sigma^{2})$.
Once the average density is found in each redshift slice, the overdensity can be calculated for each voxel in the slice from Equation 2. As above, the distribution of $\log(1+\delta_{\rm gal})$ in the slice was fit with a $2\sigma$-clipped Gaussian resulting in a $\mu_{\delta}(z)$ and $\sigma_{\delta}(z)$ for that slice (bottom left panel of Figure 4). It should be noted that, given that protostructures are extended in redshift space over multiple redshift slices, it is possible that an entire redshift slice is over- or underdense. Thus, a fifth-order polynomial was fit to $\mu_{\delta,z}$ and $\sigma_{\delta,z}$ to better characterize the general variation across redshift (see right panels of Figure 4), resulting in measurements for the parameters as a function of redshift, $\mu_{\delta,5}(z)$ and $\sigma_{\delta,5}(z)$.
With these in hand, we now have a metric for characterizing the local overdensity in each voxel of the VMC map by representing it as
@@
Figure 4: Left: Distributions (with median set to zero) of log-density (top), and local overdensity (bottom) for all of the voxels in the redshift bin centered on $z\sim 2.51$. The Gaussian and $2\sigma$ clipped Gaussian (red and green, respectively) are fit over both distributions. Right: Distribution of the mean overdensity (top) and standard deviation of overdensity (bottom) as a function of redshift. The functions are fit with a fifth-order polynomial (green line) in order to smooth over small fluctuations in redshift space.
@@ 3.2 Identifying protostructures
Due to their inherently inhomogeneous overdensity profiles, the precise spatial extent of protostructures can be difficult to define and somewhat ambiguous. Therefore, we adopted criteria intended to provide an operational definition within the VMC framework which is both physically motivated and consistent with previous works. Specifically, we defined a region to be a “protostructure” if it satisfied all three of the following criteria:
@@
    it was a contiguous set of voxels with overdensity $n_{\sigma}\geq 2.5$;
@@
    it contained at least one voxel with $n_{\sigma}\geq 4.0$, corresponding to an overdensity peak;
@@
    the voxels together enclosed a total mass $M_{\rm tot}\geq 10^{13}\,M_{\odot}$.
[…]
The $n_{\sigma}\geq 4.0$ threshold appearing in criterion (ii) sets the boundary between overdensity peaks and lower-density outskirts within Hyperion and was chosen to facilitate comparison with other studies, particularly the analysis of the Elentári proto-supercluster presented in Forrest et al. (2024b). Although Elentári also resides in the COSMOS field, it lies at $z\sim 3.3$, making a direct comparison in terms of $\delta_{\rm gal}$ unphysical. We therefore adopted an identical $n_{\sigma}$-threshold to define peaks, as this metric is more readily comparable across redshift. In our maps, $n_{\sigma}\geq 4.0$ corresponds to $\log_{10}(1+\delta_{\rm gal})\sim 0.35$ at $z\sim 2.5$.
Finally, criterion (iii) roughly corresponds to a condition imposed on the average overdensity of the protostructure and the number of voxels from which it is composed. The mass of the voxels is estimated by using (from Steidel et al. 1998)
@@
where $\bar{\rho}$ is the mean comoving density of the universe, $\langle\delta_{\rm gal}\rangle$ is the mean overdensity of the voxels defining the peak, $V$ is the combined comoving volume of the voxels, and $b$ is the bias factor, (2.55, as derived at $z\sim 2.5$ by Durkalec et al. 2015). It should be noted that, while this mass estimate may be rough, in this work it is used only as a method for differentiating between regions that are overdense because they are potentially protostructures, and regions that are overdense simply due to cosmological perturbations in density. Given the robustness of our field sample (see Figure 6) and the functional form of Equation 4, any uncertainties in the bias factor will have no meaningful effects on the conclusion of this paper.
With this framework established, Hyperion was defined as the most massive structure in the COSMOS field. By considering only voxels within Hyperion above higher $n_{\sigma}$ thresholds, we could further isolate its internal overdensity structure. Using the $n_{\sigma}\geq 4.0$ threshold appearing in criterion (ii), we defined the ”peaks” of Hyperion to be contiguous sets of voxels within Hyperion with $n_{\sigma}\geq 4.0$. The “outskirts” of Hyperion were then defined as the remaining voxels associated with Hyperion that lie in the overdensity range $2.5\leq n_{\sigma}<4.0$. A more detailed description of the structure of Hyperion is provided in Appendix A.
Finally, we note that the results of this study are not sensitive to the precise choice of overdensity thresholds used to define protostructures in the VMC maps. Following the framework presented in Section 4, we have verified that the SMF of Hyperion varies smoothly as a function of the overdensity threshold used to define it (i.e., as a function of $n_{\sigma}$ in criterion (i)), up to $n_{\sigma}\gtrsim 4.5$, beyond which the available sample size of galaxies becomes too small for robust statistical comparison.